# Array-Carrying Symbolic Execution for Function Contract Generation

### v1

@@ 1 Introduction
In program analysis, functional contract generation aims at the automated generation of contracts that specifies the input-output relationship of an individual function. The input-output relationship typically involves the pre- and post-condition, and may also include assigns information (from the ANSI/ISO C Specification Language (ACSL) assertion format) that specifies possible modifications over memory segments. The problemis fundamental in program analysis, as it facilitates interprocedural analysis by providing the basic information of functions.In this work, we focus on the function contract generation of array manipulating programs.
@@
While existing approaches tackle function contracts with array manipulations from different perspectives, they are limited in various aspects. For example, abstract interpretation has developed relational and segment-oriented domains that can reason about many array invariants [38, 14, 22, 16, 7]. However, for fully automated contract synthesis, preserving path-sensitive disjunctive loop-exit information and translating it into precise merged array-segment assigns information remains challenging. Model checking/SMT solving approaches can also infer summaries/contracts in multiple settings [32, 12, 21, 3, 19]. Nevertheless, depending on the encoding and property setup, many workflows are geared toward proving specific assertions, and may require additional summarization layers to produce reusable function-level contracts of the form $\langle\textit{Pre},\textit{Post},\textit{Assign}\rangle$. Instrumentation focuses on manual specification of ghost variables, and hence does not have a dedicated treatment of contiguous array segment. LLM approaches usually require manual guidance from experts to avoid producing incorrect results caused by hallucinations. As a result, it remains to be a challenge a framework to handle array segments and invariants in function contract generation.
@@
  • Our symbolic execution is designed to receive numerical and array invariants from external invariant generators and propagate them along the execution of program statements. This allows our framework to leverage loop invariants to generate function contracts for general properties. Moreover, the propagation of invariants also produce tight precondition for subsequent loop analysis.
@@
Experimental evaluation over a wide range of benchmarks (including realistic benchmarks from active libraries) show that our framework is capable of generating contracts that describe precise behavior of array manipulating functions, while previous approaches fail to handle many of these benchmarks.
Limitations. Our symbolic execution has limited ability to handle pointers. Currently, we can only handle simple pointer arithmetics with base and offset arising from array addresses, and could not handle recursive pointer structures such as linked lists and binary trees.
@@ 1.1 Related Works
Model checking. Model checking approaches [12, 21] often encode read/write operations of arrays by SMT rules and then solves target assertions with arrays. The most relevant works are by Tricera [3] and a SeaHorn-based verifier[20] that generate the function contracts by a transformation into CHC solving (via e.g. the state of the art CHC solver Eldarica [24]). A major disadvantage of existing model checking approaches is that they often follow top-down analysis so that the generated specification is often highly specific to the initial values of a program. Our approach circumvent this limitation by leverages external invariant generators in our symbolic execution.
@@ 3.1 Symbolic Execution Framework
In this section, we introduce our symbolic execution framework. We focus on the read and write aspects of arrays, and omit other technical aspects such as records and pointer arithmetics. A high-level program model is relegated to Appendix 0.A for the space constraint.
@@ 3.2 The Carrying of Invariants and Contiguous Array Segments
  • Segment Merging: The framework also merges segments to maintain coarse-grained array summaries. After a write operation or loop analysis, a syntactic merging strategy is applied. This procedure scans all segments sharing the same base address and identifies pairs of adjacent segments—e.g., $s_{1}=\langle(\beta_{b},\beta_{o}),\beta_{L1},\mathcal{C}_{1}\rangle$ and $s_{2}=\langle(\beta_{b},\beta_{o}+\beta_{L1}),\beta_{L2},\mathcal{C}_{2}\rangle$. If their content summaries are syntactically equivalent,
    the two are merged into a single, larger segment $s_{\mathit{new}}=\langle(\beta_{b},\beta_{o}),\beta_{L1}+\beta_{L2},\mathcal{C}_{1}\rangle$.
@@
As our symbolic execution framework relies on specific invariant generators, we describe the current state of the art on the invariant generation with numerical values and arrays. Numerical invariant generation (e.g., linear and polynomial invariants) has been extensively studied, including abstract interpretation  [16, 7], constraint solving [11, 28, 10, 41, 61, 26, 4, 39], recurrence analysis [29, 58], etc. Invariant generation with arrays are mostly tackled via contiguous array segments [14, 22, 38, 35]. Array invariants can also be solved via CHC solving [19]. A recent result investigates instrumentation [2] to aid the generation of array invariants.
@@ 3.3 Prototype Implementation Detail
Our prototype is implemented on top of the Clang/LLVM toolchain [36] that comprises about $19000$ lines in C++. We use the Clang front-end to obtain the abstract syntax tree of the input C program and traverse the AST to construct the intermediate program syntax $S$ and the initial symbolic configuration described in Section 3.1. Our framework incorporates records and pointer arithmetics. For pointer arithmetic, the current implementation supports simple forms (e.g., linear base-plus-offset patterns), but does not yet handle complex shape analysis such as linked lists or binary trees. For each analyzed function, the framework generates corresponding ACSL annotations [5] that encode the synthesized function contract, including preconditions, postconditions, and assigns clauses. These annotations are then fed to Frama-C [31], where the WP plugin is used to automatically discharge the resulting proof obligations and thereby validate the soundness of the inferred contracts with respect to the original C code. Loop reasoning in the prototype is realized via the plugin-based LoopAnalyzer introduced in Appendix 0.C.
@@ 4 Evaluation
We evaluate our symbolic execution framework by comparing it against AutoDeduct [50], a state-of-the-art toolchain capable of generating ACSL annotations for C programs. The toolchain incorporates Saida [52], which requires TriCera [19, 53, 57], and the Interface Specification Propagator (ISP) [51] to manage, propagate and emit specifications. We do not compare with popular software verification tools such as Seahorn [23] and CPAChecker [6] since they usually require assertions to verify, while ours do not require assertions and tries to generate precise contracts without manually given assertions. We also do not compare with large language models as they follow a completely different principle (see Appendix 0.D for details). All experiments were conducted in a Docker container based on Ubuntu 24.04.3 LTS running on a WSL2 host (version 6.6.87.2-microsoft-standard-WSL2) on a machine with AMD Ryzen 7 8845H CPU and 8 GB RAM [^1] We provide the source code, Docker environment, and benchmark scripts at: https://anonymous.4open.science/r/ACSLAutoGen-4753/README.md..
@@ Appendix 0.A A Basic Program Model
Formally, let the set Addr of memory addresses simply be the set of non-negative integers, and $X$ be the set of integer and array variables appearing in the underlying program. An address map is an injective map $\mathcal{A}:X\to\text{\sl Addr}$ from the set of program variables $X$ to the address space Addr that specifies the addresses of the variables, for which the injective property ensures disjoint memory allocation of variables. A memory map is a map $M:\text{\sl Addr}\to\text{Val}$ that for each address $a$, $M(a)$ is the integer element that resides at this address. It follows that given an address map $\mathcal{A}$ and a memory map $M$, the notation $M(\mathcal{A}(x))$ with an integer variable $x$ denotes the value of the variable, and the notation $M(\mathcal{A}(ar)+i)$ denotes the array entry $ar[i]$. Here, the index $i$ is also called the address offset w.r.t the address of the array variable $ar$. Since all addresses are statically assigned and disjoint, array access operation can be treated as direct access to symbolic addresses, and therefore we do not consider aliasing of different array variables.
Usually, one requires that array contents are allocated as contiguous and non-overlapping memory ranges. Formally, if $A,B\in X$ are array variables with implicit lengths $\lvert A\rvert$ and $\lvert B\rvert$ (measured in addressable cells), then their address memory intervals $\mathcal{A}(A),\,\mathcal{A}(A)+\lvert A\rvert-1]$ and $\mathcal{A}(B),\,\mathcal{A}(B)+\lvert B\rvert-1]$ should be disjoint, i.e., $I_{A}\cap I_{B}=\varnothing$. In this work, we however take a lightweight treatment of array bounds so that each array variable $ar$ is treated as simply a pointer. This allows the index $i$ in an array access $ar[i]$ unbounded. We do not eliminate array out-of-bounds explicitly.
@@
  • $T$ is a finite set of transitions. Each transition $\tau\in T$ is a tuple $\langle\ell,\ell^{\prime},g,F\rangle$ that specifies the execution of a statement moving from the location $\ell$ to the location $\ell^{\prime}$ in the underlying program, for which
    the component $g$ is the guard condition of the transition that is a formula over integer and array variables,
    and $F$ is the update function of the execution that maps the current memory map $M$ and address map $\mathcal{A}$ to the next memory map $F(M,\mathcal{A})$ after the execution of the transition.
@@
  • $\ell_{0}=\ell^{*}$ and $\sigma_{0}\models\theta$, i.e., the execution path starts from the initial location and the initial program state fulfills the initial condition $\theta$;

### v2

# arXiv:2602.23216v2[cs.PL] 27 Feb 2026

@@ 1 Introduction
In program analysis, functional contract generation aims at the automated generation of contracts that specify the input-output relationship of an individual function. The input-output relationship typically involves preconditions and postconditions, and may also include assigns information (from the ANSI/ISO C Specification Language (ACSL) assertion format) that specifies possible modifications over memory segments. The problem is fundamental in program analysis, as it facilitates interprocedural analysis by providing the basic information of functions. In this work, we focus on function contract generation for array-manipulating programs.
@@
While existing approaches tackle function contracts with array manipulations from different perspectives, they are limited in various aspects. For example, abstract interpretation has developed relational and segment-oriented domains that can reason about many array invariants [38, 14, 22, 16, 7]. However, for fully automated contract synthesis, preserving path-sensitive disjunctive loop-exit information and translating it into precise merged array-segment assigns information remains challenging. Model checking/SMT solving approaches can also infer summaries/contracts in multiple settings [32, 12, 21, 3, 19]. Nevertheless, depending on the encoding and property setup, many workflows are geared toward proving specific assertions, and may require additional summarization layers to produce reusable function-level contracts of the form $\langle\textit{Pre},\textit{Post},\textit{Assign}\rangle$. Instrumentation focuses on manual specification of ghost variables, and hence does not have a dedicated treatment of contiguous array segments. LLM approaches usually require manual guidance from experts to avoid producing incorrect results caused by hallucinations. As a result, it remains a challenge to build a framework that handles array segments and invariants in function contract generation.
@@
  • Our symbolic execution is designed to receive numerical and array invariants from external invariant generators and propagate them along the execution of program statements. This allows our framework to leverage loop invariants to generate function contracts for general properties. Moreover, the propagation of invariants also produces tight preconditions for subsequent loop analysis.
@@
Experimental evaluation over a wide range of benchmarks (including realistic benchmarks from active libraries) shows that our framework is capable of generating contracts that describe precise behavior of array-manipulating functions, while previous approaches fail to handle many of these benchmarks.
Limitations. Our symbolic execution has limited ability to handle pointers. Currently, we can only handle simple pointer arithmetic with base and offset arising from array addresses, and cannot handle recursive pointer structures such as linked lists and binary trees.
@@ 1.1 Related Works
Model checking. Model checking approaches [12, 21] often encode read/write operations of arrays by SMT rules and then solve target assertions with arrays. The most relevant works are by Tricera [3] and a SeaHorn-based verifier[20] that generate function contracts through a transformation into CHC solving (via e.g. the state-of-the-art CHC solver Eldarica [24]). A major disadvantage of existing model checking approaches is that they often follow top-down analysis, so the generated specification is often highly specific to the initial values of a program. Our approach circumvents this limitation by leveraging external invariant generators in our symbolic execution.
@@ 3.1 Symbolic Execution Framework
In this section, we introduce our symbolic execution framework. We focus on the read and write aspects of arrays, and omit other technical aspects such as records and pointer arithmetic. A high-level program model is relegated to Appendix 0.A for space constraints.
@@ 3.2 The Carrying of Invariants and Contiguous Array Segments
  • Segment Merging: The framework also merges segments to maintain coarse-grained array summaries. After a write operation or loop analysis, a syntactic merging strategy is applied. This procedure scans all segments sharing the same base address and identifies pairs of adjacent segments—e.g., $s_{1}=\langle(\beta_{b},\beta_{o}),\beta_{L1},\mathcal{C}_{1}\rangle$ and $s_{2}=\langle(\beta_{b},\beta_{o}+\beta_{L1}),\beta_{L2},\mathcal{C}_{2}\rangle$. If their content summaries are syntactically equivalent, the two are merged into a single, larger segment $s_{\mathit{new}}=\langle(\beta_{b},\beta_{o}),\beta_{L1}+\beta_{L2},\mathcal{C}_{1}\rangle$.
@@
As our symbolic execution framework relies on specific invariant generators, we describe the current state of the art on the invariant generation with numerical values and arrays. Numerical invariant generation (e.g., linear and polynomial invariants) has been extensively studied, including abstract interpretation  [16, 7], constraint solving [11, 28, 10, 41, 61, 26, 4, 39], recurrence analysis [29, 58], etc. Invariant generation with arrays is mostly tackled via contiguous array segments [14, 22, 38, 35]. Array invariants can also be solved via CHC solving [19]. A recent result investigates instrumentation [2] to aid the generation of array invariants.
@@ 3.3 Prototype Implementation Detail
Our prototype is implemented on top of the Clang/LLVM toolchain [36] that comprises about $19000$ lines in C++. We use the Clang front-end to obtain the abstract syntax tree of the input C program and traverse the AST to construct the intermediate program syntax $S$ and the initial symbolic configuration described in Section 3.1. Our framework incorporates records and pointer arithmetic. For pointer arithmetic, the current implementation supports simple forms (e.g., linear base-plus-offset patterns), but does not yet handle complex shape analysis such as linked lists or binary trees. For each analyzed function, the framework generates corresponding ACSL annotations [5] that encode the synthesized function contract, including preconditions, postconditions, and assigns clauses. These annotations are then fed to Frama-C [31], where the WP plugin is used to automatically discharge the resulting proof obligations and thereby validate the soundness of the inferred contracts with respect to the original C code. Loop reasoning in the prototype is realized via the plugin-based LoopAnalyzer introduced in Appendix 0.C.
@@ 4 Evaluation
We evaluate our symbolic execution framework by comparing it against AutoDeduct [50], a state-of-the-art toolchain capable of generating ACSL annotations for C programs. The toolchain incorporates Saida [52], which requires TriCera [19, 53, 57], and the Interface Specification Propagator (ISP) [51] to manage, propagate, and emit specifications. We do not compare with popular software verification tools such as Seahorn [23] and CPAChecker [6] since they usually require assertions to verify, while our approach does not require assertions and instead tries to generate precise contracts without manually given assertions. We also do not compare with large language models as they follow a completely different principle (see Appendix 0.D for details). All experiments were conducted in a Docker container based on Ubuntu 24.04.3 LTS running on a WSL2 host (version 6.6.87.2-microsoft-standard-WSL2) on a machine with AMD Ryzen 7 8845H CPU and 8 GB RAM [^1] We provide the source code, Docker environment, and benchmark scripts at: https://doi.org/10.5281/zenodo.18601854.
@@ Appendix 0.A A Basic Program Model
Formally, let the set Addr of memory addresses simply be the set of non-negative integers, and $X$ be the set of integer and array variables appearing in the underlying program. An address map is an injective map $\mathcal{A}:X\to\text{\sl Addr}$ from the set of program variables $X$ to the address space Addr that specifies the addresses of the variables, for which the injective property ensures disjoint memory allocation of variables. A memory map is a map $M:\text{\sl Addr}\to\text{Val}$ that for each address $a$, $M(a)$ is the integer element that resides at this address. It follows that given an address map $\mathcal{A}$ and a memory map $M$, the notation $M(\mathcal{A}(x))$ with an integer variable $x$ denotes the value of the variable, and the notation $M(\mathcal{A}(ar)+i)$ denotes the array entry $ar[i]$. Here, the index $i$ is also called the address offset w.r.t the address of the array variable $ar$. Since all addresses are statically assigned and disjoint, array access operations can be treated as direct access to symbolic addresses, and therefore we do not consider aliasing of different array variables.
Usually, one requires that array contents are allocated as contiguous and non-overlapping memory ranges. Formally, if $A,B\in X$ are array variables with implicit lengths $\lvert A\rvert$ and $\lvert B\rvert$ (measured in addressable cells), then their address memory intervals $\mathcal{A}(A),\,\mathcal{A}(A)+\lvert A\rvert-1]$ and $\mathcal{A}(B),\,\mathcal{A}(B)+\lvert B\rvert-1]$ should be disjoint, i.e., $I_{A}\cap I_{B}=\varnothing$. In this work, we take a lightweight treatment of array bounds so that each array variable $ar$ is treated as a pointer. This allows the index $i$ in an array access $ar[i]$ to be unbounded. We do not eliminate array out-of-bounds explicitly.
@@
  • $T$ is a finite set of transitions. Each transition $\tau\in T$ is a tuple $\langle\ell,\ell^{\prime},g,F\rangle$ that specifies the execution of a statement moving from the location $\ell$ to the location $\ell^{\prime}$ in the underlying program, for which the component $g$ is the guard condition of the transition that is a formula over integer and array variables, and $F$ is the update function of the execution that maps the current memory map $M$ and address map $\mathcal{A}$ to the next memory map $F(M,\mathcal{A})$ after the execution of the transition.
@@
  • $\ell_{0}=\ell^{*}$ and $\sigma_{0}\models\theta$, i.e., the execution path starts from the initial location and the initial program state fulfills the initial condition $\theta$.